The following figures are available for Insure (Pty) Ltd:

Actual (2018)

Budget (2018)

January

February

March

April

May

Sales

96 000

120 000

60 000

160 000

200 000

Purchases

100 000

50 000

160 000

84 000

60 000

Sales commission

400

600

1 200

800

400

Wages

800

800

800

?

?

  1. 40% of all sales are for cash.
  2. 35% of all purchases are for cash.
  3. Collection from debtors are as follows:
    20% is collected in the month of the sale and 2% discount is granted on these collections.
    60% is collected in the month following the month of sale.
    15% is collected in the second month following the month of sale.
    The remaining 5% is written off as a bad debt.
  4. Insure (Pty) Ltd will make an investment of R50 000 in fixed deposit on 01 April 2018.
  5. Interest of R500 per month on the fixed deposit is expected to be received from 30 April 2018.
    1. Creditors are paid 2 months after date of invoice.
    2. Rent expense amounts to R8 000 per month and is payable on the first day of each month.
    3. Wages will be increased by 10% in April and will increase by a further 5% in May 2018.
    4. A new machine is expected to be purchased for R10 000 cash in May 2018.
    5. Commission is paid to sales personnel in the month following the month in which it was earned.
    6. On 31 March 2018, Insure (Pty) Ltd had a bank overdraft of R28 000.

      Required:

Prepare the debtors‘ collection schedule for April and May 2018. (10)

Prepare the cash budget for April and May 2018.   (20)
Answer #1
Please give positive ratings so I can keep answering. If you have any queries please comment. Thanks!
Insure Pty Ltd.
Credit Sales January February March April May
Sales     96,000.00 120,000.00      60,000.00      160,000.00 200,000.00 A
Cash Sales @ 40%     38,400.00     48,000.00      24,000.00        64,000.00      80,000.00 B=A*40%
Credit Sales     57,600.00     72,000.00     36,000.00        96,000.00 120,000.00 C=A-B
Collection from Debtors January February March April May
20% in the month of the sale     11,520.00     14,400.00        7,200.00        19,200.00      24,000.00 D=C*20%
Less: 2% discount          230.40          288.00           144.00              384.00           480.00 E=D*2%
Net collected in the month of the sale     11,289.60     14,112.00        7,056.00        18,816.00      23,520.00 F=D-E
60% in month following the month of sale     34,560.00      43,200.00        21,600.00      57,600.00 G=60% of C of previous month
15% in second month following sale        8,640.00        10,800.00        5,400.00 H=60% of C of previous to previous month
Collection from Debtors        51,216.00     86,520.00 I=F+G+H
Credit Purchases January February March April May
Purchases 100,000.00     50,000.00 160,000.00        84,000.00      60,000.00 J
Cash Purchases @ 35%     35,000.00     17,500.00      56,000.00        29,400.00      21,000.00 K=J*35%
Credit Purchases     65,000.00     32,500.00 104,000.00        54,600.00     39,000.00 L=J-K
Payment to Creditors January February March April May
Credit Purchases     65,000.00     32,500.00 104,000.00        54,600.00      39,000.00 L
Creditors are paid 2 months after date of invoice      65,000.00        32,500.00 104,000.00 M= L of previous to previous month
Payment to Creditors     65,000.00        32,500.00 104,000.00 M
Payment of Wages January February March April May
Wages paid          800.00          800.00           800.00              880.00           924.00 N
10% increase in April i.e. 800+10%of 800= R 880.
5% increase in May i.e. 880+5%of 880= R 924.
Cash budget April May
Beginning Balance       (28,000.00)    (33,864.00)
Add:
Cash Sales        64,000.00      80,000.00 B
Collection from Debtors        51,216.00      86,520.00 I
Interest Income from Fixed Deposit              500.00           500.00
Cash Available        87,716.00 133,156.00
Disbursements
Fixed deposit        50,000.00                    -  
Cash Purchases        29,400.00      21,000.00 K
Payment to Creditors        32,500.00 104,000.00 M
Rent payment          8,000.00        8,000.00
Payment of Wages              880.00           924.00 N
Sales commission              800.00           400.00
Machinery purchased                       -        10,000.00
Total cash payments      121,580.00 144,324.00
Ending balance (Overdraft)      (33,864.00) (11,168.00)